# Pseudoentanglement from tensor networks

### v1

@@ Abstract
The derivation of (7) makes no reference to the MPS structure and thus proceeds unchanged. Once we have a partition function with positive Boltzmann weights
$e^{-(E^{+}[\{\sigma_{i}^{\prime},\tau_{i}\}]-E^{+}[\{e,e\}])}=\prod_{(i,j)\in{\rm edges}}\left\{\begin{aligned} &1&\text{ if }\alpha_{i}=\alpha_{j},\\ &O(1/\chi_{ij})&\text{ otherwise.}\end{aligned}\right.,$ (S15)
we can bound it above by simply dropping edges—i.e. setting the associated factor to 1 regardless of the spins $\alpha_{i,j}$. Dropping edges until we obtain a spanning tree $\mathcal{T}$ of the graph, as sketched in Fig. 2(b,d), and moving to bond variables $\beta_{ij}\equiv\alpha_{i}\alpha_{i}^{-1}$ starting from the pinned site $\sigma_{1}^{\prime}=e$ and covering all links in the spanning tree, we get
$\sum_{\{\sigma_{i}^{\prime},\tau_{i}\}\text{ n.u.}}e^{-(E^{+}[\{\sigma_{i}^{\prime},\tau_{i}\}]-E^{+}[\{e,e\}])}$ $\leq\prod_{\ell\in\mathcal{T},\text{ solid}}\left(\sum_{\beta}\chi_{\ell}^{|\beta|-m}\right)\prod_{\ell\in\mathcal{T},\text{ dashed}}\left(\sum_{\beta}\frac{|\textsf{Wg}_{d_{\ell}}(\beta)|}{\textsf{Wg}_{d_{\ell}}(e)}\right)-1$
$\leq\prod_{\ell\in\mathcal{T},\text{ solid}}[1+O(m^{2}/\chi_{\ell})]\prod_{\ell\in\mathcal{T},\text{ dashed}}[1+O(m^{2}/d_{\ell})]-1$
@@
${\rm Tr}(\rho_{A}^{2})=\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}e^{-E_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]},$ (S19)
where the energy term $E_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]$ comprises both bulk terms (overlaps or Weingarten functions between neighboring permutations) and boundary terms (overlaps between $\sigma$’s and $e$ in $A$, and between $\sigma$’s and $s$ in $\bar{A}$, respectively). Since we are interested in an upper bound on the purity (i.e. a lower bound on the entropy), we can switch to positive Boltzmann weights:
${\rm Tr}(\rho_{A}^{2})\leq\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}\left|e^{-E_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]}\right|=\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}e^{-E^{+}_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]},$ (S20)
The graph contains two types of bonds: overlaps between permutations (solid lines in Fig. 2(d)) and Weingarten functions (dashed lines in Fig. 2(d)). There are $n_{o}$ of the former type and $n_{w}$ of the latter. The overlaps contribute a factor of either $\chi^{2}$ (if the two permutations they connect are the same) or $\chi$ (if they are opposite); the Weingarten functions are either $(\chi^{6}-1)^{-1}$ or $\chi^{-1}(\chi^{6}-1)^{-1}$ respectively. We factor out a $\chi^{2}$ from each of the $n_{o}$ overlap bonds and $(\chi^{6}-1)^{-1}$ from each of the $n_{w}$ Weingarten bonds. In all, this gives a factor of $\chi^{2n_{o}}/(\chi^{6}-1)^{n_{w}}$; since $n_{w}=3n_{o}$ (from a counting of total number of legs in the network), we get
${\rm Tr}(\rho_{A}^{2})$ $\leq(1-\chi^{-6})^{n_{w}}\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}e^{-\delta E^{+}_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]}\leq[1+O(N/\chi^{6})]\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}e^{-\delta E^{+}_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]}.$ (S21)
The energy terms are now normalized in such a way that they contribute a factor of 1 if the two permutations are equal, and $1/\chi$ or $1/\chi^{3}$ respectively if the permutations are different. Since the partition function depends only on the relative value of neighboring spins, if we include (pinned) spins at the bounary ($s$ in $A$ and $e$ in $\bar{A}$) then the sum over site variables $\{\sigma_{i},\tau_{i}\}$ can be rewritten in terms cuts $\gamma$ through the bonds of the network, with
@@
and $|\gamma|$ the number of bonds cut by $\gamma$ (with the convention that cutting a Weingarten link contributes 3 bonds in this case). This is already a RT formula with a fluctuating cut. Then, owing to the large bond dimension $\chi$, fluctuations of the cut are suppressed and one obtains ‘mininum-cut’ prescription. To this end, let us (very loosely) upper bound the number of possible cuts of a given length $\ell$ as $\binom{n_{b}}{\ell}$, with $n_{b}=n_{e}+n_{o}$ the total number of bonds in the network (on hyperbolic tilings, where bulk and boundary scale in the same way, we have $n_{b}=O(N)$). Then, letting $\ell_{\rm min}=\min_{\gamma_{A}}|\gamma_{A}|$, we have

### v2

@@ Abstract
The derivation of Eq. (7) makes no reference to the MPS structure and thus proceeds unchanged. Once we have a partition function with positive Boltzmann weights
$e^{-(E^{+}[\{\sigma_{i}^{\prime},\tau_{i}\}]-E^{+}[\{e,e\}])}=\prod_{(i,j)\in{\rm edges}}\left\{\begin{aligned} &1&\text{ if }\alpha_{i}=\alpha_{j}\\ &O(1/\chi_{ij})&\text{ otherwise}\end{aligned}\right.,$ (S15)
we can bound it above by simply dropping edges, i.e., setting the associated factor to 1 regardless of the spins $\alpha_{i,j}$. We drop edges until we obtain a spanning tree $\mathcal{T}$ of the graph, as sketched in Fig. 2(b,d), and move to bond variables $\beta_{ij}\equiv\alpha_{i}\alpha_{i}^{-1}$ starting from the pinned site $\sigma_{1}^{\prime}=e$ and covering all links in the spanning tree. This gives
$\sum_{\{\sigma_{i}^{\prime},\tau_{i}\}\text{ n.u.}}e^{-(E^{+}[\{\sigma_{i}^{\prime},\tau_{i}\}]-E^{+}[\{e,e\}])}$ $\leq-1+\prod_{\ell\in\mathcal{T},\text{ solid}}\left(\sum_{\beta}\chi_{\ell}^{|\beta|-m}\right)\prod_{u\in\mathcal{T},\text{ dashed}}\left(\sum_{\beta}\frac{|\textsf{Wg}_{d_{u}}(\beta)|}{\textsf{Wg}_{d_{u}}(e)}\right)$
$\leq-1+\prod_{\ell\in\mathcal{T},\text{ solid}}[1+O(m^{2}/\chi_{\ell})]\prod_{u\in\mathcal{T},\text{ dashed}}[1+O(m^{2}/d_{u})]$
@@
${\rm Tr}(\rho_{A}^{2})=\frac{1}{D}\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}e^{-E_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]},$ (S19)
where the energy term $E_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]$ comprises both bulk terms (overlaps or Weingarten functions between neighboring permutations) and boundary terms (overlaps between $\sigma$’s and $e$ in $A$, and between $\sigma$’s and $s$ in $\bar{A}$, respectively). The prefactor of $1/D=\chi^{-L}$ comes from the normalization of the input Bell pairs, $\frac{1}{\sqrt{\chi}}\sum_{j=1}^{\chi}\ket{jj}$, contributing a factor of $\chi^{-2}$ to $\rho_{A}^{2}$ for each pair of output qudits, hence overall $(\chi^{-2})^{L/2}=\chi^{-L}=1/D$.
Since we are interested in an upper bound on the purity (i.e. a lower bound on the entropy), we can switch to positive Boltzmann weights:
${\rm Tr}(\rho_{A}^{2})\leq\frac{1}{D}\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}\left|e^{-E_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]}\right|=\frac{1}{D}\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}e^{-E^{+}_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]},$ (S20)
The graph contains two types of bonds: overlaps between permutations (solid lines in Fig. 2(d)) and Weingarten functions (dashed lines in Fig. 2(d)). The overlaps contribute a factor of either $\chi^{2}$ (if the two permutations they connect are the same) or $\chi$ (if they are opposite); the Weingarten functions are either $(\chi^{6}-1)^{-1}$ or $\chi^{-3}(\chi^{6}-1)^{-1}$ respectively. Let us define $n_{o}$ and $n_{w}$ as the number of overlap bonds and Weingarten bonds, respectively; by counting the number of edges in the graph, we have $n_{o}=3n_{w}+L/2$ (recall $L$ is the number of output qudits). Factoring out a $\chi^{2}$ from each of the $n_{o}$ overlap bonds and $(\chi^{6}-1)^{-1}$ from each of the $n_{w}$ Weingarten bonds, we get a factor of $\chi^{2n_{o}}/(\chi^{6}-1)^{n_{w}}=\chi^{L}(1-\chi^{-6})^{-n_{w}}$, thus
${\rm Tr}(\rho_{A}^{2})$ $\leq(1-\chi^{-6})^{-n_{w}}\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}e^{-\delta E^{+}_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]}\leq[1+O(N/\chi^{6})]\sum_{\{\sigma_{i},\tau_{i}\in S_{2}\}}e^{-\delta E^{+}_{A,\bar{A}}[\{\sigma_{i},\tau_{i}\}]}.$ (S21)
The energy terms are now normalized in such a way that they contribute a factor of 1 if the two permutations are equal, and $1/\chi$ (for overlap bonds) or $1/\chi^{3}$ (for Weingarten bonds) respectively if the permutations are different. Since the partition function depends only on the relative value of neighboring spins, the pinned values of the spins at the boundary ($s$ in $A$ and $e$ in $\bar{A}$) allow us to turn the sum over site variables $\{\sigma_{i},\tau_{i}\}$ into a sum over bond variables. This in turn can be rewritten in terms of cuts $\gamma$ through the bonds of the network (a bond is ‘cut’ if and only if its spin variable is the transposition $s\in S_{2}$), giving
@@
where $|\gamma|$ is the number of bonds intersected by $\gamma$ (with the convention that a Weingarten bond counts as 3 bonds in this case). This is a RT formula with a fluctuating cut; then, owing to the large bond dimension $\chi$, fluctuations of the cut are suppressed and one obtains ‘mininum-cut’ prescription.
To show this, let us (very loosely) upper bound the number of possible cuts of a given length $\ell$ as $\binom{n_{b}}{\ell}$, with $n_{b}=n_{w}+n_{o}$ the total number of bonds in the network (on hyperbolic tilings, where bulk and boundary scale in the same way, we have $n_{b}=O(N)$). Then, letting $\ell_{\rm min}=\min_{\gamma_{A}}|\gamma_{A}|$, we have